Convert a speed of 4.60 mi/s

4.60 mi/s to units of kilometers per minute. Show the unit analysis by dragging the conversion factors into the unit‑factor slots.


1
Expert's answer
2021-08-30T15:05:22-0400

One mile is "1mi \approx 1.61km" of kilometers. One second is "1s = 1\/60 min". Thus, obtain:


"4.6\dfrac{mi}{s} = 4.6\cdot \dfrac{1.61km}{1/60min} = 4.6\cdot 96.6\dfrac{km}{min} \approx 444 \dfrac{km}{min}"

Answer. "444 \dfrac{km}{min}".